OpenStudy (anonymous):

\[\frac{(a^4-b^2)}{a-b}\] is equal to ?

OpenStudy (anonymous):

\[\frac{(a^2-b)(a^2+b)}{a-b}\]

OpenStudy (anonymous):

its equal to something, but it is more elegant this way.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

\[\frac{(a^2)^2-(b)^2)}{(a-b)}\] \[\frac{(a^2+b)(a^2-b)}{a-b}\] after then ?

OpenStudy (anonymous):

thats it ......... it is cannot simplified

OpenStudy (anonymous):

options : 1)a^2+b^2 2) (a+b)(a^2-b^2) 3) (a+b)(a^2+b^2) 4) none of these

OpenStudy (anonymous):

options are now given help me now please

OpenStudy (anonymous):

i think the ans is 4

OpenStudy (anonymous):

iam sure!!!!

OpenStudy (anonymous):

how ?